# ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT

### v1

@@ Abstract.
Technical question and answering (Q&A) sites such as Stack Overflow have become an important source for software developers to seek knowledge. However, code snippets on Q&A sites are usually uncompilable and semantically incomplete for compilation due to unresolved types and missing dependent libraries, which raises the obstacle for users to reuse or analyze Q&A code snippets. Prior approaches either are not designed for synthesizing compilable code or suffer from a low compilation success rate. To address this problem, we propose ZS4C, a lightweight approach to perform zero-shot synthesis of compilable code from incomplete code snippets using Large Language Model (LLM). ZS4C operates in two stages. In the first stage, ZS4C utilizes an LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template. In the second stage, ZS4C fixes compilation errors caused by incorrect import statements and syntax errors through collaborative work between ChatGPT and a compiler. We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO against the SOTA approach SnR. Compared with SnR, ZS4C improves the compilation rate from 63% to 87.6%, with a 39.3% improvement. On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1.
imcomplete code snippets, Large Language Model, Program synthesis, API inference, Prompt Engineering, ChatGPT
@@ 1. Introduction
However, the majority of the code snippets on Q&A sites are uncompilable, which raises the obstacle for users to reuse or analyze these code snippets (Terragni et al., 2016; Rubei et al., 2020). Those code snippets are usually uncompilable and semantically incomplete for compilation due to unresolved types (e.g., what is the fully qualified name (FQN) for a type DateFormat) and missing dependent libraries (e.g., which library to import for DataFormat) (Terragni et al., 2016; Mishne et al., 2012; Dong et al., 2022). Although missing such implementation details could be synthesized manually, it is tedious and often requires substantial familiarity with various libraries from developers. Furthermore, such manual synthesis effort is, however, not scalable for software engineering research based on Q&A sites’ code snippets (Zhang et al., 2021, 2019; Gao et al., 2015; Mishne et al., 2012). A technique to automatically repair code snippets by resolving the missing types and dependent libraries for compilation will significantly save developers time and facilitate the research based on such code snippets.
@@
In this study, we introduce ZS4C, an approach designed to perform zero-shot synthesis of compilable code from incomplete code snippets using LLM. ZS4C operates in two stages. In the first stage, ZS4C utilizes LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template. We apply the widely adopted self-consistency method to improve the robustness of the inference. In the second stage, ZS4C employs the Multi-Agent Conversation framework (Wu et al., 2023) to fix compilation errors caused by incorrect import statements and syntax errors. Specifically, ZS4C uses a compiler as an agent to provide feedback, i.e., compilation error logs, which serve as additional context for ChatGPT to correct the errors. Compared with prior approaches, ZS4C is lightweight and does not require any training data, any pre-defined rules or heuristics, or program analysis like constructing AST.
We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO consisting of 267 code snippets manually collected from Stack Overflow posts (Phan et al., 2018) against the SOTA approach SnR. ZS4C improves the compilation rate of SnR, from 63% to 87.6%, with a 39.3% improvement. On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1 score. Our ablation analysis shows that both components in ZS4C play a crucial role in the effectiveness of ZS4C. For instance, by employing the component conversation between a compiler and ChatGPT, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement. Our parameter analysis reveals that ZS4C is not sensitive to parameters. Lastly, we conducted a failure analysis and our results sheds light on the limitations of using LLMs for synthesizing code, such as hallucination (e.g., fake FQN, and partially correct FQN) and unexpected code modification from LLMs.
@@
  • We propose ZS4C, a novel zero-shot LLM-based approach to automatically synthesize compilable code for code snippets, which employs the multi-agent conversation framework to boost the successful rate of synthesizing compilable code for code snippets.
  • We conducted extensive experiments to demonstrate the effectiveness of ZS4C compared with SOTA approach, and other baselines that mimic how developers use ChatGPT for resolving compilation issues for incomplete code snippets.
@@
  • We made our code and dataset publicly available (dat, 2023) to encourage future research on studying partial code synthesis and related approaches.
@@ 2.1. Automated Synthesis of Compilable Code from Stack Overflow
Another work from Dong et al (Dong et al., 2022), in which they proposed a tool named SnR, that can automatically infer FQNs, recommend libraries, and create import statements for code snippets, eventually making code compilable. SnR is a constraint-based approach. It first analyzes the incomplete code snippet and extracts constraints that capture the relationships between different elements in the code. Next, SnR utilizes a knowledge base built from a collection of real Java libraries which contains information about classes, methods, and dependencies. The extracted constraints are then solved using a Datalog solver called Soufflé (Dong et al., 2022). This solver uses the knowledge base to find solutions that satisfy the constraints. Once the constraints are solved, SnR transforms the solutions into import statements. These import statements are added to the code snippet to resolve missing dependencies. They have compared their tool with the state-of-the-art FQN inference approach named COSTER (Saifullah et al., 2019). Experiments on a benchmark of 267 code snippets from Stack Overflow show that SnR correctly infers 91.0% of the import statements, which makes 73.8% of the snippets compile, compared to 36.0% of the import statements and 9.0% of the snippets by COSTER (Saifullah et al., 2019). Given these results, we consider SnR as our baseline approach.
One of the key tasks in the automated synthesis of compilable code snippets is to identify the fully qualified name (FQN) for APIs (FQN inference(Huang et al., 2022). Specifically, identifying the types of API elements in incomplete code can help identify the dependent libraries. Existing studies addressed FQN inference in different ways: Baker extracts constraints from code snippets and uses a naive constraint solving algorithm to infer FQNs (Subramanian et al., 2014); StatType formulates the task as a phrase-based machine translation task - translating the code snippet into a target sentence that contains the FQNs of the API elements (Phan et al., 2018) and builds a statistical model from existing compilable source code; COSTER (Saifullah et al., 2019) is another statistical model-based approach, it calculates the likelihood of appearing context tokens and the FQN of each API element, and then ranks the FQNs based on three different scores: likelihood, context similarity, and name similarity scores; Huang et al. (Huang et al., 2022) treats FQN inference as clozy-style fill-in-blank language task. They utilized the pre-trained CodeBERT model (Feng et al., 2020) and fine-tuned it using prompt learning with FQN-annotated code examples, sampled from target libraries. The prompt learning stimulates the MLM to recognize FQN syntax and usage patterns. In our approach, we do not directly perform FQN inference, instead, we skip it and directly focus on import statement inference.
@@ 2.2. Using LLMs for Software Engineering Tasks
With the breakthrough of LLMs in the field of natural language processing, there’s a burgeoning interest in investigating the potential of LLMs to enhance software engineering tasks. Researchers have pre-trained LLMs on publicly available code, resulting in the development of specialized code LLMs like CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), Codex (Chen et al., 2021), and CodeGen (Nijkamp et al., 2022). In addition to these pre-trained code models, another research trajectory focuses on optimizing the application of LLMs to software engineering tasks such as code generation (Liu et al., [n. d.]), program repair (Wei et al., 2023), testing (Schäfer et al., 2023), and beyond. Given the remarkable capability of LLMs to learn from zero or minimal examples, there’s a growing exploration into how in-context learning and prompt engineering can elevate LLM-based applications in software engineering. For instance, Gao et al. (Gao et al., 2023) investigated the influence of the selection, sequence, and quantity of demonstration examples on in-context learning performance in code intelligence tasks, namely, code summarization, bug fixing, and program synthesis. They found that meticulously curated demonstrations can significantly outperform conventional demonstration construction techniques. Different from the aforementioned research, we focus on designing a new prompt engineering method to synthesize compilable code for incomplete code snippets on Q&A sites by leveraging LLM as a knowledge base. We introduce ZS4C, a lightweight zero-shot approach that decomposes the task into two stages, and enables a collaborative conversation between the LLM and a compiler to resolve compilation errors.
@@ 3. Methodology
We present the overview of our approach ZS4C in Figure 1. ZS4C is comprised of two components: Import Statement Inference and Conversational Error Fixing. In Import Statement Inference, we leverage a large language model (LLM) to identify the missing dependent libraries (by identifying missing import statements) in the code snippet. In Conversational Error Fixing, we further fix compilation errors caused by the wrong dependent libraries inferred by Import Statement Inference and other compilation errors (e.g., syntax error) to boost the compilation success rate of a code snippet, through the conversation between a compiler and the LLM. For simplicity, in the following section, we refer to Import Statement Inference as ImportStInfer and Conversational Error Fixing as ConversationFixing.
@@ 3.1.1. Prompt design
#Instruction Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Do not check for any import statements in the code. Only give correct imports by not using wildcard imports. Please note that you need to pay close attention and your response should be specific and accurate. Avoid repetition and must not generate wrong and nonexistent imports:” #Query {$input\_code\_snippet$}
@@
We design the instructions to guide and control the output of the model for our task step by step in Main_Instruction. We first ask the model not to check the code for any import statements if present. [^2] Note that, in line with previous studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), the input code snippets utilized for evaluation in this study do not contain any import statements. This design directs the model’s attention exclusively to the body of the code rather than any existing import statements, which could be incorrect. Moreover, we observed that the model sometimes generates responses including “no imports are found”, likely stemming from a misinterpretation of the task. To mitigate this issue, we explicitly instruct the model to avoid seeking import statements from the input code snippet. Next, we instruct the model to provide only correct imports and to avoid using wildcard imports. Wildcard import refers to the mechanism by which one can import all classes and interfaces from a particular package without specifying their fully qualified names by using the asterisk (*) character. For instance, in Java, one can use “import java.util.*;” to import all classes from the “java.util” package, including the ArrayList, HashMap, and Scanner classes, among others, rather than individually importing classes, e.g., “import java.util.ArrayList”. We instruct the model to avoid wildcard imports for several reasons: 1) wildcard imports may cause naming conflicts if multiple classes with the same name exist in different packages; 2) we want to steer the model towards identifying the exact classes required rather than resorting to the short-cut wildcard imports; 3) wildcard imports are generally discouraged in practice because it can compromise the code readability and introduce unintended side-effects that can be very difficult to debug (Van Oort et al., 2021). Following the recommended best practice for prompt design mentioned in  (White et al., 2023), we add a sentence “Please note that you need to pay close attention and your response should be specific and accurate.” to explicitly emphasize the importance of specificity and accuracy for the LLM to prioritize. We also observed that in some cases, the model generates duplicate import statements and incorrect import statements that referenced non-existent libraries. To mitigate this issue, we added a sentence “Avoid repetition and must not generate wrong and nonexistent imports” to the prompt template.
@@ 3.1.2. Self-consistency
To obtain reliable and coherent results from the model, we employ the self-consistency method proposed by Wang et al. (Wang et al., 2022). As indicated by its name, self-consistency simply asks a model to run the same prompt multiple times and determines the final answer based on the principle of frequency majority. More specifically, we run the same prompt for $K$ times and the model generates $K$ lists of import statements for our query. We count the frequency of each list and take the list with the highest frequency as the final result. If at least two lists have the same frequency, we repeat the process (i.e., running the same prompt for another $K$ times) until we break the tie. For instance, we query the model using the same prompt 5 times, and the model generates five lists, $L_{1}=\{A,B,C\}$, $L_{2}=\{B,A,C\}$, $L_{3}=\{A,B,C,D\}$, $L_{4}=\{A,B\}$, and $L_{5}=\{C,B,A\}$, we take $\{A,B,C\}$ as the final result as it (the combination of three libraries - A, B, C) appears 3 times with the highest frequency. In this study, we empirically set $K$ to 10, which achieves the best performance (see results of RQ4).

### v3

# MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=

@@ Abstract.
Technical Q&A sites are valuable for software developers seeking knowledge, but the code snippets they provide are often uncompilable and incomplete due to unresolved types and missing libraries. This poses a challenge for users who wish to reuse or analyze these snippets. Existing methods either do not focus on creating compilable code or have low success rates. To address this, we propose ZS4C, a lightweight approach for zero-shot synthesis of compilable code from incomplete snippets using Large Language Models (LLMs). ZS4C operates in two stages: first, it uses an LLM, like GPT-3.5, to identify missing import statements in a snippet; second, it collaborates with a validator (e.g., compiler) to fix compilation errors caused by incorrect imports and syntax issues. We evaluated ZS4C on the StatType-SO benchmark and a new dataset, Python-SO, which includes 539 Python snippets from Stack Overflow across the 20 most popular Python libraries. ZS4C significantly outperforms existing methods, improving the compilation rate from 63% to 95.1% compared to the state-of-the-art SnR, marking a 50.1% improvement. On average, ZS4C can infer more accurate import statements (with an F1 score of 0.98) than SnR, with an improvement of 8.5% in the F1.
Incomplete code snippets, Large Language Model, Program synthesis, API inference, Prompt engineering, ChatGPT
@@ 1. Introduction
However, the majority of the code snippets on Q&A sites are uncompilable, which raises the obstacle for users to reuse or analyze these code snippets (Terragni et al., 2016; Rubei et al., 2020). Those code snippets are usually uncompilable and semantically incomplete for compilation due to unresolved types (e.g., what is the fully qualified name (FQN) for the type DateFormat?) and missing dependent libraries (e.g., which library to import for DateFormat?) (Terragni et al., 2016; Mishne et al., 2012; Dong et al., 2022). Although missing such implementation details could be synthesized manually, it is tedious and often requires substantial familiarity with various libraries from developers. Furthermore, such manual synthesis effort is, however, not scalable for software engineering research based on Q&A sites’ code snippets (Zhang et al., 2021, 2019; Gao et al., 2015; Mishne et al., 2012). Besides, Integrated Development Environments (IDEs) sometimes fail to provide accurate import statements for the code snippets (JetBrains Community, 2018). This issue is also prevalent in LLM-generated code snippets, which can occasionally miss the appropriate import statements for successful execution (Pan et al., 2024). A technique to automatically repair code snippets by resolving the missing types and dependent libraries for compilation will significantly save developers time and facilitate the research based on such code snippets.
@@
In this study, we introduce ZS4C, an approach designed to perform zero-shot synthesis of compilable code from incomplete code snippets using LLM. ZS4C operates in two stages. In the first stage, ZS4C utilizes LLM (e.g., ChatGPT), to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template. We apply the widely adopted self-consistency method to improve the robustness of the inference. In the second stage, ZS4C employs iterative conversations between an LLM and a validator (e.g., compiler) to fix errors caused by incorrect import statements and syntax errors. Specifically, ZS4C uses a validator to provide feedback, e.g., compilation error logs, which serve as additional context for an LLM to correct the errors. Compared with prior approaches, ZS4C is lightweight and does not require any training data, any pre-defined rules or heuristics, or program analysis like constructing AST.
We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO consisting of 267 code snippets manually collected from Stack Overflow posts (Phan et al., 2018) and a new dataset called Python-SO that contains 539 Python code snippets collected from Stack Overflow across 20 popular Python libraries. ZS4C outperforms baselines on all studied datasets. We evaluated ZS4C with two models, GPT-3.5 and GPT-4. ZS4C improves the compilation rate of the SOTA approach SnR, from 63% to 95.1%, with a 50.1% improvement. On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 8.5% in the F1 score. Our ablation analysis shows that both components in ZS4C play a crucial role in the effectiveness of ZS4C. For instance, by employing the component conversation between a validator and GPT-3.5, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement. Our parameter analysis reveals that ZS4C is not sensitive to parameters. Lastly, we conducted a failure analysis and our results shed light on the limitations of using LLMs for synthesizing code, such as hallucination (e.g., fake FQN, and partially correct FQN) and unexpected code modification from LLMs.
@@
  • We propose ZS4C, a novel zero-shot LLM-based approach to automatically synthesize compilable code for code snippets, which employs iterative conversations LLM and validator (e.g., compiler) to boost the successful rate of synthesizing compilable code for code snippets.
  • We conducted extensive experiments to demonstrate the effectiveness of ZS4C compared with the SOTA approach, and other baselines that mimic how developers use LLMs (e.g., ChatGPT) for resolving compilation issues for incomplete code snippets.
@@
  • We construct new data for Python libraries. We made our code and dataset publicly available (dat, 2023) to encourage future research on studying partial code synthesis and related approaches.
@@ 2.1. Automated Synthesis of Compilable Code from Stack Overflow
Another work from Dong et al (Dong et al., 2022), in which they proposed a tool named SnR, that can automatically infer FQNs, recommend libraries, and create import statements for code snippets, eventually making code compilable.
SnR is a constraint-based approach. It first analyzes the incomplete code snippet and extracts constraints that capture the relationships between different elements in the code. Next, SnR utilizes a knowledge base built from a collection of real Java libraries which contains information about classes, methods, and dependencies. The extracted constraints are then solved using a Datalog solver called Soufflé (Dong et al., 2022). This solver uses the knowledge base to find solutions that satisfy the constraints. Once the constraints are solved, SnR transforms the solutions into import statements. These import statements are added to the code snippet to resolve missing dependencies. They have compared their tool with the state-of-the-art FQN inference approach named COSTER (Saifullah et al., 2019). Experiments on a benchmark of 267 code snippets from Stack Overflow show that SnR correctly infers 91.0% of the import statements, which makes 73.8% of the snippets compile, compared to 36.0% of the import statements and 9.0% of the snippets by COSTER (Saifullah et al., 2019). Given these results, we consider SnR as our baseline approach.
One of the key tasks in the automated synthesis of compilable code snippets is to identify the fully qualified name (FQN) for APIs (FQN inference(Huang et al., 2022). Specifically, identifying the types of API elements in incomplete code can help identify the dependent libraries.
Existing studies addressed FQN inference in different ways: Baker extracts constraints from code snippets and uses a naive constraint solving algorithm to infer FQNs (Subramanian et al., 2014); StatType formulates the task as a phrase-based machine translation task - translating the code snippet into a target sentence that contains the FQNs of the API elements (Phan et al., 2018) and builds a statistical model from existing compilable source code; COSTER (Saifullah et al., 2019) is another statistical model-based approach, it calculates the likelihood of appearing context tokens and the FQN of each API element, and then ranks the FQNs based on three different scores: likelihood, context similarity, and name similarity scores; Huang et al. (Huang et al., 2022) treats FQN inference as clozy-style fill-in-blank language task. They utilized the pre-trained CodeBERT model (Feng et al., 2020) and fine-tuned it using prompt learning with FQN-annotated code examples, sampled from target libraries. The prompt learning stimulates the MLM to recognize FQN syntax and usage patterns. In our approach, we do not directly perform FQN inference, instead, we skip it and directly focus on import statement inference.
@@ 2.2. Using LLMs for Software Engineering Tasks
With the breakthrough of LLMs in the field of natural language processing, there’s a burgeoning interest in investigating the potential of LLMs to enhance software engineering tasks. Researchers have pre-trained LLMs on publicly available code, resulting in the development of specialized code LLMs like CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), Codex (Chen et al., 2021), and CodeGen (Nijkamp et al., 2022). In addition to these pre-trained code models, another research trajectory focuses on optimizing the application of LLMs to software engineering tasks such as code generation (Liu et al., [n. d.]), program repair (Wei et al., 2023; Liu et al., 2024a), testing (Schäfer et al., 2023), and beyond. Given the remarkable capability of LLMs to learn from zero or minimal examples, there’s a growing exploration into how in-context learning and prompt engineering can elevate LLM-based applications in software engineering. For instance, Gao et al. (Gao et al., 2023) investigated the influence of the selection, sequence, and quantity of demonstration examples on in-context learning performance in code intelligence tasks, namely, code summarization, bug fixing, and program synthesis. They found that meticulously curated demonstrations can significantly outperform conventional demonstration construction techniques. Liu et al. introduced the LLM-CompDroid framework, which combines the strengths of LLMs and traditional tools for repairing configuration compatibility bugs in Android apps (Liu et al., 2024a). Their experimental results demonstrate a significant enhancement in bug resolution performance and surpassing the state-of-the-art tool. Different from the aforementioned research, we focus on designing a new prompt engineering method to synthesize compilable code for incomplete code snippets on Q&A sites by leveraging LLM as a knowledge base and collaborating with a compiler. We introduce ZS4C, a lightweight zero-shot approach that decomposes the task into two stages, and enables a collaborative conversation between the LLM and a compiler to resolve compilation errors.
@@ 3. Methodology
We present the overview of our approach ZS4C in Figure 1. ZS4C is comprised of two components: Import Statement Inference and Conversational Error Fixing. In Import Statement Inference, we leverage a large language model (LLM) to identify the missing dependent libraries (by identifying missing import statements) in the code snippet. In Conversational Error Fixing, we further fix compilation errors caused by the wrong dependent libraries inferred by Import Statement Inference and other compilation errors (e.g., syntax error) to boost the compilation success rate of a code snippet, through the conversation between a validator such as a compiler which could provide error log, and the LLM. For simplicity, in the following section, we refer to Import Statement Inference as ImportStInfer and Conversational Error Fixing as ConversationFixing.
@@ 3.1.1. Prompt design
Prompt Template - 1 for Import Statement Inference #Instruction
Presentation_Guide: “Reply with to-the-point answer, no elaboration.”
Main_Instruction (for Java): “Do not check for any import statements in the code. Only give correct imports by not using wildcard imports. Please note that you need to pay close attention and your response should be specific and accurate. Avoid repetition and must not generate wrong and nonexistent imports:”
Main_Instruction (for Python): “Only give correct import statements for the attached code. Please note that you need to pay close attention and your response should be specific and accurate. Avoid repetition and must not generate wrong imports:”
#Query
{$input\_code\_snippet$}
@@
We design the instructions to guide and control the output of the model for our task step by step in Main_Instruction. We first ask the model to ignore any import statements in the given input code if present. [^2] Note that, in line with previous studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), the input code snippets utilized for evaluation in this study do not contain any import statements. However, code snippets on SO sometimes come with import statements. This design directs the model’s attention exclusively to the body of the code rather than any existing import statements, which could be incorrect. Moreover, we observed that the model sometimes generates responses including “no imports are found”, likely stemming from a misinterpretation of the task (e.g., to find import statements in the provided code snippet). To mitigate this issue, we explicitly instruct the model to avoid seeking import statements from the input code snippet. Another issue is wildcard import, which refers to the mechanism by which one can import all classes and interfaces from a particular package without specifying their fully qualified names by using the asterisk (*) character. For instance, in Java, one can use “import java.util.*;” to import all classes from the “java.util” package, including the ArrayList, HashMap, and Scanner classes, among others, rather than individually importing classes, e.g., “import java.util.ArrayList”. Wildcard may cause several issues: 1) wildcard imports may cause naming conflicts if multiple classes with the same name exist in different packages; 2) we want to steer the model towards identifying the exact classes required rather than resorting to the short-cut wildcard imports; 3) wildcard imports are generally discouraged in practice because it can compromise the code readability and introduce unintended side-effects that can be very difficult to debug (Van Oort et al., 2021). Therefore, we instruct the model to provide only correct imports and to avoid using wildcard imports. Following the recommended best practice for prompt design mentioned in  (White et al., 2023), we add a sentence “Please note that you need to pay close attention and your response should be specific and accurate.” to explicitly emphasize the importance of specificity and accuracy for the LLM to prioritize. Without this sentence, LLM sometimes outputs explanations with the inferred import statements and brings us challenges in extracting the import statement from the output. We also observed that in some cases, the model generates duplicate import statements and incorrect import statements that reference non-existent libraries (e.g., hallucination). To mitigate this issue, we added the sentence “Avoid repetition and must not generate wrong and nonexistent imports” to the prompt template.
@@ 3.1.2. Self-consistency
To obtain reliable and coherent results from the model, we employ the self-consistency method proposed by Wang et al. (Wang et al., 2022). As indicated by its name, self-consistency simply asks a model to run the same prompt multiple times and determines the final answer based on the principle of frequency majority. More specifically, we run the same prompt for $K$ times and the model generates $K$ lists of import statements for our query. We count the frequency of each list and take the list with the highest frequency as the final result. If at least two lists have the same frequency, we repeat the process (i.e., running the same prompt for another $K$ times) until we break the tie. For instance, if we query the model using the same prompt 5 times, and the model generates five lists, $L_{1}=\{A,B,C\}$, $L_{2}=\{B,A,C\}$, $L_{3}=\{A,B,C,D\}$, $L_{4}=\{A,B\}$, and $L_{5}=\{C,B,A\}$. We take $\{A,B,C\}$ as the final result as it (the combination of three libraries - A, B, C) appears 3 times with the highest frequency. In this study, we empirically set $K$ to 10, which achieves the best performance (see results of RQ3). Note that we only apply self-consistency in the Import Statement Inference component.